# Multi-finger Manipulation via Trajectory Optimization with Differentiable Rolling and Geometric Constraints

### v1

@@ V-B Baselines and Ablations
We compare our method with the following 3 methods:
@@
Planning to Contact (PtC): after pregrasping, the robot fingers are in contact with the object. Given the initial contacts, the fingertip (end effector) positions $\hat{\mathbf{p}}_{i}$ in the object frame are recorded. By linearly interpolating the object pose between the current and goal poses, we get the desired object pose for each time step. Assuming the contact points do not change in the object frame, we determine the desired fingertip locations at each time step in the world frame. An inverse kinematics (IK) problem of reaching the desired fingertip position is then solved to obtain the robot joint angles. The desired pose at the next time step is executed and the process above is iteratively repeated at each step. This method does not consider rolling or object geometry. The idea of searching the end effector pose first and then solving an IK is used in many robot manipulation tasks, especially in the motion planning domain [38, 39, 40].
@@
Model Predictive Path Integral (MPPI): MPPI is a popular model-based trajectory optimization method. We implement MPPI as proposed in [7]. We use the simulator as the dynamics model, i.e., we roll out the trajectory using the simulator. The simulator requires both configurations and their velocities as the state, while all the other methods are only given the configurations. To ensure fairness, we set the velocities input into the dynamics model to 0 due to the quasi-static assumption. We use 200 particles for sampling, i.e., at each time step, we will spawn 200 copies of the environment, each executing different action sequences sampled. Additionally, we set the sampling horizon $T=2$. We tried setting $T$ the same as that in our method, but the robot almost always dropped the object. This is because our task is sensitive and it is nearly impossible to constantly satisfy the stringent contact constraints in a lower-dimensional manifold over an extended horizon through random action sampling, resulting in dropping the object frequently and high costs for almost every trajectory. Using a longer horizon will lead to a poorer performance in our experiments.
Ablation without Geometry-Based Constraints: To demonstrate the importance of differentiable geometry information and rolling constraints, we remove the contact constraints (Eq. 11) and the kinematics constraints (Eq. 12). Instead, we replace those constraints with simplified contact constraints. Similar to the PtC method, we assume the fingertip locations in the object frame do not change:
$f_{FK,i}(\mathbf{q}_{t})-f_{trans}(\mathbf{o}_{t},\hat{\mathbf{p}}_{i})=0,$ (23)
@@
where $f_{FK,i}$ is the forward kinematics function to get the fingertip location for the $i$th finger, $f_{trans}$ returns the desired fingertip location in the world frame based on the current object pose $\mathbf{o}_{t}$. Unlike PtC, this method considers friction, joint limits, and force balances. Additionally, the optimizer selects the next object pose rather than linear interpolation.
@@ V-C Simulation experiments
Peg Alignment Screwdriver Turning
distance valid distance valid
Ours $11.64^{\circ}\pm 1.74^{\circ}$ 1.00 $\mathbf{1.72^{\circ}\pm 0.83^{\circ}}$ 1.00
PtC $31.56^{\circ}\pm 7.98^{\circ}$ 0.0 $145.50^{\circ}\pm 31.87^{\circ}$ 0.0
MPPI $\mathbf{8.62^{\circ}\pm 4.93^{\circ}}$ 0.65 $96.93^{\circ}\pm 51.75^{\circ}$ 0.0
Ablation $39.46^{\circ}\pm 3.41^{\circ}$ 0.25 $43.94^{\circ}\pm 20.72^{\circ}$ 0.0
TABLE I: Results are evaluated for 20 episodes in the simulation.
We ran 20 trials for each method on each task. The results are shown in Fig. 5 and Table I. Our method is the only one to achieve 100% validity rate and can reorient the object close to the desired orientation. Rolling behaviors are also observable in our experiments.
@@
For the PtC method, the desired end effector positions are usually unreachable, causing the IK solver to fail most of the time. For example, in the screwdriver turning task, the thumb is almost fully extended and has a limited configuration space, which requires rolling the thumb to turn the screwdriver further. Without a valid IK solution, the robot’s behavior becomes unpredictable, leading to dropping the object frequently.
The MPPI method is very effective in exploring the desired turning behavior. In the peg alignment task, it can even reorient the peg very close to the desired configurations. However, these results are somewhat misleading, since the validity rate is only 0.65, mostly because the robot drops the peg. Similarly, in the screwdriver turning task, the robot can approach the goal in the first half of a trial but fails to maintain stability, leading to frequently dropping the screwdriver in the second half. The dropping results in a very large distance to the goal. Dexterous manipulation tasks can be very sensitive to the control input, e.g., dropping can occur if some finger fails to make the desired contact. MPPI’s reliance on random action sampling makes it challenging to sample action sequences that can hold the object stably in the hand, contributing to high variance in the results.
@@
The ablation method, which excludes geometry-based constraints, still considers the rest of the constraints including force balance and friction. Those constraints are designed to ensure that the robot will not drop the object. However, simply tracking the desired fingertip location can lead to unmodeled rolling behavior, pushing the object away from the goal (see the screwdriver task in Fig. 5), or preventing the desired rolling behavior from occurring (see the peg alignment task in Fig. 5). The ablation results highlight the importance of modeling rolling in dexterous manipulation.
Computation time is important for whether the method can be applied in an MPC framework. For our method, excluding the warm-up step, the average computation time for each step is 8.23s and 8.92s for peg alignment and screwdriver turning, respectively. The warm-up time is 65.18s and 55.25s. For PtC, it is 34.23s and 31.57s, since we use sufficiently long iterations for solving IK. For MPPI, it is 9.44s and 11.46s, due to a large number of particles used. For the ablation, it is 6.52s and 9.11s, excluding the warm-up time. The warm-up time is 65.85s and 56.93s.
@@
Clearly, none of these methods have adequate computation time for real-time MPC. However, it is encouraging that our method’s computation time is better than baselines’ and comparable to the ablation. We did not focus on computational efficiency for this work and all methods were implemented in Python. We expect that optimizing the code and implementing our approach in C++ would lead to significant speed improvements.
V-D Real-World Experiment
@@
In the real-world experiment, we place a screw in a loose slot as the Allegro hand cannot exert large torques for tightening the screw. Our screwdriver is 3D printed since the allegro hand is larger than a human hand and the size of a normal screwdriver is not suitable for this robot. Similarly to a regular screwdriver, a high-friction tape is attached to the screwdriver body. We use Aruco tags [41] to estimate the poses of the screwdriver and the robot hand (see Fig. 6).
@@
We ran 10 trials of the real-world experiment. The qualitative results are shown in Fig. 6. The robot drops the screwdriver twice. In the remaining 8 trials of not dropping it, the average distance to the goal is 22.4 degrees. The sim2real gap can be attributed to: (1) Perception error: inaccurate object pose estimation (2) Imperfect joint modeling: unmodeled factors such as backlash and joint friction. The hardware experiment shows that our method is not only effective in simulation but also transferable to the real world, though perception and modeling challenges remain.
Figure 6: Physical screwdriver turning experiment
@@ VI CONCLUSION
Notably in our experiments, tasks such as peg alignment and screwdriver turning have been shown to benefit significantly from the differentiable formulations of contacts and finger rolling. By adopting a sampling method to approximate the finger geometries and integrating it with the object SDF, we estimate the contact-related variables in a differentiable manner. The formulations are used by a gradient-based trajectory optimizer to handle non-trivial geometries and extend the capabilities of dexterous manipulation. In future work we would like to integrate our method with contact mode planning.

### v2

@@ V-B Baselines and Ablations
We compare our method with the following 4 methods:
@@
Planning to Contact (PtC): The robot starts in contact with the object. Given the initial contacts, the fingertip (end effector) positions $\hat{\mathbf{p}}_{i}$ in the object frame are recorded. By linearly interpolating the object pose between the current and goal poses, we get the desired object pose for each time step. Assuming the fingertip positions do not change in the object frame, we determine the desired fingertip locations at each time step in the world frame. An inverse kinematics (IK) problem is then solved to compute the robot’s joint angles required to reach these positions. This method does not consider rolling or object geometry. The idea of searching the end effector pose first and then solving an IK is used in many robot manipulation tasks, especially in the motion planning domain [43, 44, 45].
@@
Model Predictive Path Integral (MPPI): MPPI is a popular model-based trajectory optimization method. We implement MPPI as proposed in [7], using the simulator as the dynamics model to roll our trajectories, similar to [46]. Since the simulator requires both configurations and velocities as the state, we provide MPPI with additional velocity information, whereas other methods only utilize configurations. For valve turning, we use 50 particles for sampling, i.e., at each time step, we spawn 50 copies of the environment, each executing different action sequences sampled. The sampling horizon is set to be 4. We experimented with extending the horizon, but the performance showed no significant improvement. For the remaining tasks, we use 500 particles. Additionally, we set the sampling horizon $T=2$. We attempted setting $T$ the same as that in our method, but the robot almost always dropped the object due to the sensitive nature of the tasks: it is nearly impossible to constantly satisfy the stringent contact constraints in a lower-dimensional manifold over an extended horizon through random action sampling, resulting in dropping the object frequently and high costs for almost every trajectory. Longer horizons generally result in poorer performance in our experiments. Since MPPI performs rollouts and evaluates within the exact same environment, it should, in theory, reach the goal for all tasks given enough sampling budget and an appropriate cost function. To ensure a fair comparison, we match the sampling budget to the runtime of our method.
Reinforcement Learning: Although RL has a different problem setup , requiring training on the specific task prior to evaluation, we include it because it is a popular approach for multi-finger manipulation. We use HORA [41], as it also addressed in-hand reorientation, though with the hand facing upwards. While HORA introduces a method for sim-to-real transfer, we evaluate it solely in simulation experiments, eliminating the need for transfer. For each task, we train three different seeds and train until it has fully converged.
@@
Ablation without Geometry-Based Constraints: To demonstrate the importance of differentiable geometry information and rolling constraints, we remove the contact constraints (Eq. 8) and the kinematics constraints (Eq. 9). Instead, we replace those constraints with simplified contact constraints. Similar to the PtC method, we assume the fingertip locations in the object frame do not change: $f_{FK,i}(\mathbf{q}_{t})-f_{trans}(\mathbf{o}_{t},\hat{\mathbf{p}}_{i})=0,$ where $f_{FK,i}$ is the forward kinematics function to get the fingertip location for the $i$th finger, $f_{trans}$ returns the desired fingertip location in the world frame based on the current object pose $\mathbf{o}_{t}$. Unlike PtC, this method considers friction, joint limits, and force balances. Additionally, the optimizer selects the next object pose rather than linear interpolation.
@@ V-C Simulation experiments
We ran 30 trials for each method on each task. The results are shown in Table I. Our method is the only one to achieve validity rates close to 1 across all tasks while coming close to the desired orientation. Rolling behaviors are also observable in our experiments (see video). Our method does sometimes produce invalid trajectories. This is due to either an insufficient optimization budget or constraint violation between time steps. Specifically, our method only considers constraint at the discretized time steps. However, these constraints may be violated when the robot executes actions and transitions between neighboring time steps.
@@
For the PtC method, the desired end effector positions are usually unreachable, causing the IK solver to fail frequently. For example, in the screwdriver turning task, the thumb is almost fully extended and has a limited configuration space, which requires rolling the thumb to turn the screwdriver further. Without a valid IK solution, the robot’s behavior becomes unpredictable, leading to dropping the object frequently. However, it achieves a validity rate of $1.0$ on the cuboid turning task, as the thumb and middle require only minimal movement and these two fingers provide additional stability to prevent object dropping.
The MPPI method is very effective in exploring the desired turning behavior. It can achieve a similar distance to the goal as our method, though with a slightly lower validity rate. It even has slightly better performance on the valve turning and cuboid alignment task, as it does not need to adhere to the assumption of a fixed contact mode and quasi-static execution. However, MPPI does not consider the robustness explicitly and tends to end up with sensitive trajectories. When executing MPPI in the real-world setup, since the dynamics of the simulator no longer match the evaluation environment, i.e., the real-world setup, the performance is much worse. We will discuss it more in Sec. V-D.
In our experiment, the RL method faces significant challenges with exploration: It is heavily penalized for dropping the object in the reward function, and it’s hard to sample a long sequence of actions that adheres to the stringent contact required to prevent object dropping. As a result, it often converges to a conservative policy. For instance, in the cuboid alignment task, the RL method places the cuboid on the gray box to avoid dropping it, completely abandoning the objective of reorienting to the goal. However, we do not claim that RL methods are inherently unsuitable for our benchmark. With appropriate reward engineering and curriculum learning, it is possible to train an RL for in-hand reorientation with the hand facing downwards [1]. Our goal is to highlight the non-triviality of training a successful RL agent for our benchmark.
@@
The ablation method, which excludes geometry-based constraints, still considers the rest of the constraints to prevent dropping the object. However, simply tracking the desired fingertip location can lead to unmodeled rolling behavior, pushing the object away from the goal, as demonstrated in the screwdriver task, or preventing the desired rolling behavior from occurring as demonstrated in the cuboid alignment task. In general, tasks that inherently require more finger rolling exhibit a larger performance gap between the ablation and our method, highlighting the importance of modeling rolling in dexterous manipulation.
Task Ours Ablation PtC MPPI RL
w.u. online w.u. online online w.u. online online
VT 12.9 s 3.9 s 17.8 s 5.3 s 8.7 s 13.5 s 4.5 s $5\times 10^{-4}$ s
ST 66.5 s 10.0 s 65.8 s 9.8 s 28.3 s 78.0 s 16.4 s $6\times 10^{-4}$ s
CA 64.9 s 10.1 s 70.5 s 11.0 s 46.4 s 67.0 s 14.4 s $5\times 10^{-4}$ s
CT 39.9 s 5.6 s 46.5 s 7.0 s 8.7 s 66.1 s 13.3 s $4\times 10^{-4}$ s
CR 39.8 s 5.8 s 43.4 s 6.4 s 50.7 s 54.3 s 28.9 s $5\times 10^{-4}$ s
TABLE II: Computation time for the experiments. w.u. means warm up.
@@
The computation time is shown in Tab. II. For PtC, long iterations for solving IK result in extended runtimes when no valid solution is found. Clearly, none of these methods except RL have adequate computation time for real-time MPC. We did not focus on computational efficiency for this work and all methods were implemented in Python. We expect that optimizing the code and implementing our approach in C++ will lead to significant speed improvements. In the hardware experiment in Sec. V-D, we reduce online iterations to speed up execution with minimal impact on performance.
V-D Real-World Experiments
@@
We consider the screwdriver turning and cuboid alignment for real-world experiments. For screwdriver turning, we place a screw in a loose slot as the Allegro hand cannot exert large torques for tightening the screw. Our screwdriver is 3D-printed because the Allegro hand is larger than a human hand, making a standard screwdriver unsuitable for this robot. Similarly to a regular screwdriver, a high-friction tape is attached to the screwdriver body. The high-friction tape is also used in the cuboid alignment task. We use Aruco tags [47] to estimate the poses of the screwdriver, the cuboid, and the robot hand (see Fig. 1). The PID values are manually adjusted to closely approximate a critically damped response.
Task Method Distance to Goal Validity
Screwdriver Turning Ours $25.02^{\circ}\pm 6.66^{\circ}$ 0.9
MPPI No valid trial 0.0
Cuboid Alignment Ours $13.28^{\circ}\pm 1.35^{\circ}$ 0.9
MPPI No valid trial 0.0
TABLE III: Hardware experiment results for 10 trials
We compare our method with MPPI, the best-performing baseline over 10 trials. The results are shown in Tab. III. For our method, the robot drops the object once in both tasks. The average distance to the goal is very similar to the results in the simulation. MPPI plans trajectories without considering constraints, making these trajectories sensitive to errors. As a result, perception noise and the dynamics mismatch between simulation and the real world lead to MPPI having no valid trials.
@@
The sim2real gap can be attributed to: (1) Perception errors: inaccurate object pose estimation, and (2) imperfect joint modeling: unmodeled factors such as backlash and joint friction. Hardware experiments show our method is effective in simulation and transferable to the real world, though perception and modeling challenges remain.
@@ VI CONCLUSION
Experiments on our proposed benchmark suggest that sensitive multi-finger manipulation tasks benefit significantly from differentiable contacts and finger-rolling formulations. By adopting a sampling method to approximate the finger geometries and integrating it with the object SDF, we estimate the contact-related variables differentiably. The formulations enable a gradient-based trajectory optimizer to handle non-trivial geometries and extend the capabilities of dexterous manipulation. Our future work will integrate contact mode planning to allow changing finger contacts during tasks.